# SoK: Runtime Integrity

### v2

@@ 2.3 CFI & CFA: Definitions & Threat models
Control Flow Integrity (CFI)
Originally proposed by Abadi et al. [15], CFI is a policy-based mitigation against control-flow hijacking restricting the execution path of a program at runtime based on a pre-computed CFG. In principle, enforcing CFI on a target program involves:
@@
Instrumentation of the target’s indirect branch instructions with runtime checks, so-called Inline Reference Monitors (IRMs), which enforce the control flow to be compliant with CFG.
CFG can be generated statically (as proposed originally [15]) or dynamically, as seen in following proposals [81, 82]. When CFGCFG, it is generally difficult for an adversary to manipulate the control flow to alter the intended semantics of the target program without detection by the instrumentation checks. However, statically determining strict CFGs for complex programs remains an open challenge [83], leading many pratical approaches to over-approximate CFG.
Control Flow Attestation (CFA)
@@
A CFA instance starts with sending a request containing a cryptographic challenge to . Upon receiving the request, must execute the operation requested by (either specified implicitly or explicitly within the request). During the execution of the requested task, an RoT in must ensure that an authenticated log (CFLog) containing a representation of the control flow path executed during the operation is built. After execution completes, the RoT computes an authenticated integrity measurement (e.g., using a MAC or signature) over the received challenge, CFLog, and the application’s binary to produce a response token (Rep). Finally, transmits Rep to along CFLog. Given the need to securely store the secret used to authenticate Rep even when is potentially compromised, RoT implementations typically involve some form of secure hardware support.
@@
Upon receiving Rep, can use this evidence to determine if executed the expected correctly through a valid control flow path. Further, when Rep shows an invalid path, can analyze the anomalous evidence to determine its cause and potentially remediate it.
@@
Existing CFA techniques (see Section 3) use either (1) binary instrumentation along with Trusted Execution Environment (TEE) support; or (2) custom hardware modifications to generate CFLog by detecting and saving each branch destination to a dedicated and protected memory region. For techniques that use binary instrumentation, a pre-processing phase modifies the binary so that branch instructions are prepended with additional calls to a TEE-protected trusted code. Once called, the trusted code appends CFLog the current branch destination. In hardware-based techniques, custom hardware interfaces with the CPU to detect branches and record their destinations to protected memory.
CFI/CFA Coverage
@@
Granularity: The granularity of CFI/CFA mechanisms refers to the detail in which a particular control flow transfer is monitored/checked. In this work, we categorize the granularity of a specific technique as either coarse-grained or fine-grained. Since CFI and CFA have different security goals (local detection/prevention vs. providing remote evidence enabling remote detection), their granularity pertains to different aspects.
@@
A coarse-grained approach refers to broadly applied checks that are independent of specific control flow transfers within the code. In the case of CFI, this involves techniques applied based on instruction type and agnostic to individual transfers. For instance, the following CFI policies can be classified as coarse-grained: enforcing landing pads for calls/returns, checking function type/parameter for indirect calls, and restricting indirect control flow targets within the bounds of a specific sandbox/address space. Since these policies are generally applied to all control flow transfers within a specific scope and do not account for the specific details of each transfer, they are considered coarse-grained. In CFA, a scheme is deemed coarse-grained if it does not record all control flow transfers within the attested application onto CFLog.
@@
A fine-grained technique refers to mechanisms that apply a specific check or action for each control flow instruction. In the case of CFI, this entails schemes that verify each indirect target against a unique set of valid locations, rather than applying a broader rule based on the instruction type. For instance, enforcement through techniques like shadow stacks, jump-tables, or definition sets determined by data-flow analysis are considered fine-grained solutions. A CFA scheme is classified as fine-grained if it records all control flow transfers within the attested application.
@@
Sensitivity: Although closely related to granularity, the sensitivity of a certain technique describes a different characteristic. It refers to the extent to which execution context is considered for determining the set of valid targets. In this work, we categorize schemes as insensitive, context-sensitive, and path-sensitive.
@@
Techniques have insensitive enforcement if they do not consider the calling context or current execution path when defining the set of valid targets for a particular control flow transfer. As such, most coarse-grained CFI are also insensitive because they are applied based on the instruction type or sandbox/address range, ignoring the current execution path or the calling context.
@@
Context-sensitive approaches consider the calling context to determine the set of valid targets. Examples include target bounds being within a particular function/sandbox. Additionally, when a function is called at multiple locations within a second function, a context-sensitive approach might determine returns to any call site within the second function as valid. For forward edges, a context-sensitive approach allows any valid definitions that can reach the function containing the forward edge.
@@
Path-sensitive approaches determine the set of valid targets by considering both the calling context and the current executing path. For instance, shadow stacks are regarded as path-sensitive enforcement mechanisms for return addresses because they limit a return to a single call site, rather than a set of call sites irrespective of the current path. Furthermore, schemes that employ data-flow analysis techniques, such as reaching definitions or points-to analysis, to determine the valid destinations of indirect calls are considered path-sensitive.
@@
In CFI, sensitivity affects the local decision on whether a transfer constitutes a violation. In CFA, sensitivity reflects the type of analysis/detection that can be performed by based on the received evidence.
CFI/CFA Threat Models & Assumptions
@@
The security of most CFI techniques depends on the presence of added instrumentation used to enforce CFI checks. In many cases, this is attained via WX permissions for memory accesses, as shown in Figure 3. While sensible for user-space code, privileged code can typically disable WX enforcement. Therefore, most CFI approaches that target privileged code (e.g., Kernel) rule out code injection from their threat model.
@@
CFA mechanisms require an RoT to implement their attestation functionality, including acquisition and signing of relevant evidence. The RoT function can also attest the executed binary (and any instrumentation therein) as performed by regular RA. This removes the need for WX enforcement, as long as code is attested in a temporally consistent manner, i.e., code remains the same in the interim between its measurement and execution. This also makes CFA useful to verify privileged code and code that runs on single-privilege micro-controllers.
@@
Similar to other TEE-based security services, TEE-based CFA (e.g., [34, 71, 84]) assumes that any applications outside the (hardware-protected) trusted realm of the TEE (e.g., outside the Secure World in TrustZone) can be modified/compromised whereas the RoT implementation within the Secure World is trusted. This is typically supported by a secure boot of the trusted code and implicitly assumes a minimal and vulnerability-free RoT implementation. Some CFA methods, based entirely on custom hardware [36, 72, 85], eliminate the need to trust a software TCB within the TEE by implementing the CFA RoT entirely in hardware.
@@
Generally, both CFI and CFA consider underlying hardware to be trusted, focusing on software-based exploits.
@@ 3 Design Space
Figure 5 illustrates distinguishing factors in CFI and CFA, highlighting the consequences of design choices to their effectiveness and susceptibility to attack vectors. Table I presents a classification of recent work in CFI and CFA, capturing design principles of each mechanism and assessing their trade-offs. Aside from aspects related to Security Goals (defined in Section 2.3), the rest of this section elaborates on design factors. Next, Section 4 discusses the consequences of these design choices.
@@ 3.1 Different Objectives
CFI mechanisms primarily focus on locally detecting control-flow violations during the dispatching stage to prevent execution of exploited code from continuing, as depicted in in Figure 3. Here, the wording “prevention” is not to be confused with development stage memory safety defenses (see Section 1.3) that attempt to eliminate vulnerabilities altogether. In other words, CFI does not remove root-cause vulnerabilities. Instead, it impedes certain attack stages, increasing adversaries’ difficulty in achieving arbitrary code execution.
Conversely, CFA is concerned with providing execution evidence that can be verified and inspected remotely. In this case, attack detection occurs at a relatively late stage, but provides essential insights on attack behavior that can be used to respond to attacks that have evaded prevention measures. This also includes logical bugs (i.e., those not caused by a memory vulnerability) in a program’s control flow that CFI would not treat as an exception. In contrast, CFI does not aim to inform or convince a remote party of execution integrity, handling exceptions and faults locally.
When incorporated atop CFA, runtime auditing [38] aims to reliably deliver evidence to , even when a compromised attempts not to follow the CFA protocol (see Section 5), refusing to send reports to in an attempt to hide the exploit behavior.
@@ 3.2 Action Mechanisms
Action mechanisms fall into (i) enforcement, (ii) monitoring techniques, or (iii) hybrid (i.e. a combination thereof) and can be implemented in software or hardware-assisted. Early CFI designs relied heavily on enforcement using software-based instrumentation to introduce IRMs using generic instructions [86, 87, 19, 88, 89, 90, 91, 92, 82, 93, 94, 95]. More recent proposals leverage hardware extensions for specialized CFI instructions as IRMs [96, 21, 22, 26, 97, 28, 23, 98, 27].
@@
Hybrid CFI approaches use hardware features to locally save sequences of control flow transfers for delayed checking as a batch. For instance, PathArmor [104] leverages the Intel Branch Record (LBR) registers to enable implicit monitoring of execution paths, whereas PittyPAT [105] and CFI [81] mainly depend on the Intel Processor Tracing (PT) technology [111] to explicitly monitor and protect the execution integrity at runtime. SHERLOC [112] uses ARM Macro Trace Buffer (MTB) and TrustZone for delayed verification of CFI checks as a batch.
@@
CFA monitors the execution flow recording transfers to be reported in some form to a . C-FLAT [34] was the earliest CFA and used software instrumentation to insert IRMs to redirect each control flow transfer to a secure software routine housed within TrustZone. This routine extends branch destinations into a hash-chain before resuming the attested execution (and performing the branch). TinyCFA [35] shows an instrumentation approach to achieve CFA atop the lightweight Proof of Execution (PoX) architecture APEX [113] reducing hardware costs. Additionally, the work of Papamartzivanos et al. [114] utilizes Intel PT technology [111] for generating the runtime traces. LO-FAT [36] and ATRIUM [85] eliminate instrumentation requirements from C-FLAT by implementing custom hardware modules to detect control flow transfers and extend the hash-chain. While these early approaches produce evidence that minimizes storage/transmission costs (to the size of one hash digest), they result in loss of information, requiring to use the received hash digest to derive the exact control flow path for inspection. The complexity of this task grows exponentially leading to the well-known path explosion problem [115, 116].
@@
To ease verification and inspection of CFA evidence, more recent techniques [72, 71, 38, 117, 37, 84] generate CFLog as a lossless trace containing all relevant control flow evidence. For instance, OAT [71], ARI [37], and ISC-FLAT [84] leverage TEEs to securely update and store the runtime evidence. LiteHAX [72] and ACFA [38] utilize custom hardware for recording a verbatim trace.
3.3 Execution Environment
@@
Requirements in terms of the execution environments for CFI/CFA can be distinguished as: hardware-agnostic, extension-specific, and RoT-based.
@@
CFI focuses mainly on the first two types. Hardware-agnostic CFI mechanisms, e.g., LLVM-CFI [19] and Microsoft Control Flow Guard (MS-CFG) [118] can cover a variety of targets. However, portability can come at the price of performance and security guarantees. Hence, several CFI approaches use specific architectural (hardware) extensions in their local environments. Some involve custom hardware extensions specifically designed to support CFI, while others are repurposed from their other goals and integrated as building blocks into CFI. Examples of the former include Intel CET [22] and ARM Pointer Authentication [21], along with the body of work built upon them [23, 24, 26, 97, 28, 98]. Examples of the latter category include CFI using Intel PT [81] and ARM Trace Macrocell (TMC) [119].
@@
As discussed in Section 2.3, CFA necessitates RoT support to generate (and sign) remotely verifiable evidence. RoTs in current CFA are implemented via TEEs [34, 71, 84] or custom hardware changes on [36, 72, 38, 85, 35, 120]. Current CFA proposals aimed at user-space programs [117, 121] either trust the operating system kernel (the code integrity of which can be verified using static RA as supported by commodity TPMs) or rely on enclaved execution TEEs [122].

### v3

@@ 2.3 CFI & CFA: Definitions & Threat models
2.3.1 Control Flow Integrity (CFI)
Originally proposed by Abadi et al. [14], CFI is a policy-based mitigation against control flow hijacking attacks, restricting the execution path of a program at runtime based on a pre-computed CFG. In principle, enforcing CFI on a target program involves:
@@ 2.3.1 Control Flow Integrity (CFI)
Enforcement of control flow to comply with CFG through Reference Monitors (RMs), which are software- or hardware-based runtime checks that verify the target of any indirect branch instruction at runtime.
CFG can be generated statically (as proposed originally [14]) or dynamically, as seen in following proposals [84, 85]. When CFG CFG, it is generally difficult for an adversary to manipulate control flow and alter a program’s intended behavior without detection by the activated or inserted RMs. However, statically determining strict CFGs for complex programs remains an open challenge [86], leading many practical approaches to over-approximate CFG.
2.3.2 Control Flow Attestation (CFA)
@@
A CFA instance starts with sending a request containing a cryptographic challenge to . Upon receiving the request, must execute the operation requested by (either specified implicitly or explicitly within the request). During the execution of the requested task, an RoT in must ensure that an authenticated log (CFLog) containing a representation of the control flow path executed during the operation is built. After execution completes, the RoT computes an authenticated integrity measurement (e.g., using a Message Authentication Code (MAC) or signature) over the received challenge, CFLog, and the executed binary to produce a response token (CFReport). Finally, transmits CFReport to along with CFLog. Given the need to securely store the secret used to authenticate CFReport even when is potentially compromised, RoT implementations typically involve some form of secure hardware support.
@@
Upon receiving CFReport, can use this evidence to determine if executed the expected software correctly through a valid control flow path. Further, when CFReport shows an invalid path, can analyze the anomalous evidence to determine its cause and potentially remediate it.
@@
Existing CFA techniques (see Section 3) use either (1) binary instrumentation along with Trusted Execution Environment (TEE) support; or (2) custom hardware modifications to generate CFLog by detecting and saving each branch destination to a dedicated and protected memory region. For techniques that use binary instrumentation, a pre-processing phase modifies the binary so that branch instructions are prepended with additional calls to a TEE-protected trusted code. Once called, the trusted code appends to CFLog the current branch destination. In hardware-based techniques, custom hardware interfaces with the CPU to detect branches and record their destinations in protected memory.
2.3.3 CFI/CFA Coverage
We define the coverage of CFI/CFA proposals in terms of granularity and sensitivity.
@@ 2.3.3 CFI/CFA Coverage
Granularity: The granularity of CFI/CFA mechanisms refers to the detail in which a particular control flow transfer is monitored/checked. In this work, we categorize the granularity of a specific technique as either coarse-grained or fine-grained. Since CFI and CFA have different security goals (local detection/prevention vs. providing runtime evidence to a remote party), their granularity pertains to different aspects.
@@
A coarse-grained approach refers to broadly applied checks that are independent of specific control flow transfers within the code. In the case of CFI, this involves techniques applied based on instruction type and agnostic to individual transfers. For instance, the following CFI policies can be classified as coarse-grained: enforcing landing pads for calls/returns, checking function type/parameter for indirect calls, and restricting indirect control flow targets within the bounds of a specific sandbox/address space. Since these policies are generally applied to all control flow transfers within a specific scope and do not account for the specific details of each transfer, they are considered coarse-grained. In CFA, a scheme is deemed coarse-grained if it does not record all control flow transfers within the attested application into CFLog.
@@
A fine-grained technique refers to mechanisms that apply a specific check or action for each control flow instruction. In the case of CFI, this entails schemes that verify each indirect target against a unique set of valid locations rather than applying a broader rule based on the instruction type. For instance, enforcement through techniques like shadow stacks, jump-tables, or definition sets determined by data-flow analysis are considered fine-grained solutions. A CFA scheme is classified as fine-grained if it records all control flow transfers within the attested application.
@@
Sensitivity: Although closely related to granularity, the sensitivity of a certain technique describes a different characteristic. It refers to the extent to which execution context is considered for determining the set of valid targets. In this work, we categorize schemes as insensitive, context-sensitive, or path-sensitive.
@@
Techniques have insensitive enforcement if they do not consider the calling context or current execution path when defining the set of valid targets for a particular control flow transfer. As such, the majority of coarse-grained CFI mechanisms are regarded insensitive because they employ generic rules, e.g., based on the instruction type, ignoring the current execution path or the calling context.
@@
Context-sensitive approaches consider the calling context to determine the set of valid targets. Examples include target bounds being within a particular function/sandbox. Additionally, when a function is called at multiple locations within a second function, a context-sensitive approach might determine returns to any call site within the second function as valid. For forward edges, a context-sensitive approach allows any valid definition that can reach the function containing the forward edge.
@@
Path-sensitive approaches determine the set of valid targets by considering both the calling context and the current executing path. For instance, shadow stacks are regarded as path-sensitive enforcement mechanisms for return addresses because they limit a return to a single call site. Furthermore, schemes that employ data flow analysis, such as reaching definitions or points-to analysis, to determine the valid destinations of indirect branches are considered path-sensitive.
@@
In CFI, sensitivity affects the local decision on whether a transfer constitutes a violation, whereas in CFA, sensitivity reflects the type of analysis/detection that can be performed by based on the received evidence.
2.3.4 CFI/CFA Threat Models & Assumptions
@@
The security of most CFI techniques depends on the presence of added instrumentation used to enforce CFI checks. In many cases, this is attained via WX permissions for memory accesses, as shown in Figure 3. While sensible for user-space code, privileged code can typically disable WX enforcement. Therefore, most CFI approaches that target privileged code (e.g., Kernel) rule out code injection/modification from their threat model.
@@
CFA mechanisms require an RoT to implement their attestation functionality, including the acquisition and signing of relevant evidence. The RoT function can also attest the executed binary (and any instrumentation therein) as performed by regular RA. This removes the need for WX enforcement, as long as code is attested in a temporally consistent manner, i.e., code remains the same in the interim between its measurement and execution. This also makes CFA useful to verify privileged code and code that runs on single-privilege Micro Controller Units (MCUs).
@@
Similar to other TEE-based security services, TEE-based CFA (e.g., [33, 38, 43]) assumes that any application outside the (hardware-protected) trusted realm of the TEE (e.g., outside the Secure World in TrustZone) can be modified/compromised whereas the RoT implementation within the Secure World is trusted. This is typically supported by a secure boot of the trusted code and implicitly assumes a minimal and vulnerability-free RoT implementation, as vulnerabilities in the RoT can lead to full system compromise [87]. Some CFA methods eliminate the need to trust a software TCB within the TEE by implementing the CFA RoT entirely in hardware [35, 37, 36].
@@
Generally, both CFI and CFA consider the underlying hardware to be trusted, focusing on software-based exploits.
@@ 3 Design Space
Figure 5 illustrates the distinguishing factors in CFI and CFA, highlighting the consequences of design choices on their effectiveness and susceptibility to attack vectors. Accordingly, Table I presents a classification of recent work in CFI and CFA, capturing design principles of each mechanism and assessing their trade-offs. Aside from aspects related to security goals (defined in Section 2.3), the rest of this section elaborates on design factors. Afterward, Section 4 discusses the consequences of these design choices.
@@ 3.1 Different Objectives
CFI mechanisms primarily focus on locally detecting control flow violations during the dispatching stage to prevent execution of exploited code from continuing, as depicted in in Figure 3. In this context, “prevent” is not to be confused with the goal of memory safety defenses as mechanisms that aim to remove vulnerabilities (see Section 1.3). In other words, CFI does not remove root-cause vulnerabilities. Instead, it impedes certain attack stages, increasing adversaries’ difficulty in achieving arbitrary code execution.
Conversely, CFA is concerned with providing authentic execution evidence that can be verified and inspected remotely. In this case, attack detection occurs at a relatively late stage but provides essential insights into attack behavior that can be used to respond to attacks that have evaded prevention measures. This also includes logical bugs (i.e., those not caused by a memory vulnerability) in a program’s control flow that CFI would not treat as an exception. In contrast, CFI does not aim to inform or convince a remote party of execution integrity, handling exceptions and faults locally.
When incorporated atop CFA, runtime auditing [41, 42] aims to reliably deliver evidence to , even when a compromised attempts not to follow the CFA protocol (see Section 5), refusing to send reports to in an attempt to hide the exploit behavior.
@@ 3.2 Action Mechanisms
Action mechanisms fall into (i) enforcement, (ii) monitoring techniques, or (iii) hybrid (i.e. a combination thereof) and can be hardware-assisted or implemented in software. Early CFI designs relied heavily on enforcement through software-based instrumentation (SWI) using generic instructions, so-called Inline Reference Monitors (IRMs) [88, 89, 18, 90, 91, 92, 93, 94, 85, 95, 96, 97]. More recent proposals leverage hardware extensions for specialized CFI instructions as IRMs [98, 20, 21, 25, 99, 27, 22, 100, 26].
@@
Hybrid CFI approaches implement Hardware Reference Monitors (HRMs) using hardware features to locally save sequences of control flow transfers for asynchronous verification by a separate trusted software module. For instance, PathArmor [106] leverages the Intel Branch Record (LBR) registers to enable implicit monitoring of execution paths, whereas PittyPAT [107] and CFI [84] mainly depend on the Intel Processor Tracing (PT) technology [113] to explicitly monitor and verify the execution integrity at runtime. SHERLOC [114] uses ARM Micro Trace Buffer (MTB) and TrustZone for asynchronous detection of CFI violations.
@@
CFA monitors the execution flow, recording transfers to be reported in some form to a . C-FLAT [33] was the first CFA and used software instrumentation to insert IRMs, which redirect each control flow transfer to a secure software routine housed within TrustZone. This routine extends branch destinations into a hash-chain before resuming the attested execution (and performing the branch). TinyCFA [34] shows an instrumentation-based approach to achieve CFA atop a Proof of Execution (PoX) architecture, called APEX [115]. Additionally, the work of Papamartzivanos et al. [116] utilizes Intel PT technology [113] for generating the runtime traces. LO-FAT [35] and ATRIUM [36] eliminate instrumentation requirements from C-FLAT by implementing custom hardware modules to detect control flow transfers and extend the hash-chain.
While these early approaches produce evidence that minimizes storage/transmission costs (to the size of one hash digest), they result in loss of information, requiring to use the received hash digest to derive the exact control flow path for inspection. The complexity of this task grows exponentially, leading to the well-known path explosion problem [117, 118].
@@
To ease verification and inspection of CFA evidence, more recent techniques [37, 38, 41, 119, 40, 43] generate CFLog as a lossless trace containing all relevant control flow evidence. For instance, OAT [38], ARI [40], and TRACES [42] leverage TEEs to securely update and store the runtime evidence. LiteHAX [37] and ACFA [41] utilize custom hardware for recording a verbatim trace.
While CFA techniques are primarily focused on monitoring control flow events passively, recent techniques have proposed hybrid action mechanisms that monitor the control flow path while providing some enforcement capabilities. For instance, ISC-FLAT [43] creates a TEE-protected dispatcher to ensure that external system interrupts cannot stealthily modify the control flow of the application being attested. CFA+[120] leverages ARMv8.5-A’s landing pad instructions[98] in combination with selective software instrumentation to enforce a specific CFI policy and enable lightweight monitoring of the execution state, which is maintained in two reserved registers.
3.3 System Models & Execution Environments
@@
CFA and CFI concepts can apply to both high-end systems and low-end embedded devices. Additionally, the requirements of execution environments within the target platform for CFI/CFA can be distinguished as: hardware-agnostic, extension-specific, and RoT-based.
CFI focuses mainly on the first two types of execution environments, and it has primarily been applied in high-end systems. Many earlier CFI approaches are hardware agnostic and utilize SWI to monitor control flow events and locally detect violations. These CFI mechanisms, e.g., LLVM-CFI [18] and Microsoft Control Flow Guard (MS-CFG) [121], are easily portable and can cover a variety of high-end targets regardless of the particular CPU type. As they are hardware-agnostic and instrumentation-based, their system models require the presence of a Memory Management Unit (MMU) or an Operating System (OS) to maintain the integrity of the instrumentation. When applied to bare-metal embedded systems that lack MMUs, a Memory Protection Unit (MPU) is used for similar guarantees [111, 122].
@@
The portability of CFI in hardware-agnostic environments can come at the price of performance. Hence, several CFI approaches use specific architectural (hardware) extensions in their local environments. Some involve custom hardware extensions specifically designed to support CFI, while others are repurposed from their other goals and integrated as building blocks into CFI. Examples of the former include Intel CET [21] and ARM Pointer Authentication [20], along with the body of work built upon them [22, 23, 25, 99, 27, 100]. Examples of the latter category include CFI using Intel PT [84] and ARM Trace Macrocell (TMC) [123]. As extension-specific execution environments are available on both high-end and embedded platforms, CFI proposals using them can also be applied to both classes of target devices.
@@
Most existing CFA approaches target embedded platforms, including those with low-power single-core MCUs (e.g., Atmel AVR ATMega, TI MSP430) with 1-16MHz CPUs, 8-16 bit instruction set architectures (ISAs), used to execute “bare-metal” software and commonly implement IoT sensors/actuators and control systems. Additionally, CFA has been proposed for 32-bit embedded platforms with TEE support (e.g., ARM Cortex-M MCUs) that execute real-time applications. As discussed in Section 2.3, CFA necessitates RoT support to generate (and sign) remotely verifiable evidence. RoTs in CFA for embedded devices are implemented via TEEs [33, 38, 43, 42] or custom hardware changes on [35, 37, 41, 36, 34, 124]. Current CFA proposals aimed at user-space programs [119, 125, 120] either trust the OS (the code integrity of which can be verified using static RA as supported by commodity TPMs) or rely on enclaved execution TEEs [126].